Sales (13,200 units × $30 per unit) $ 396,000

Variable expenses 198,000

Contribution margin 198,000

Fixed expenses 220,500

Net operating loss $ (22,500 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,700 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $86,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $33,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.50 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,900?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $50,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,600 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,600)?
Answer #1
Req 1 :
Contribution margin per unit = Contribution margin / Units sold = 198000 / 13200 15
CM ratio = Contribution margin / Sales = 198000 / 396000 50%
Break-even point in units = Fixed expenses / Contribution margin per unit = 220500 / 15 14700
Break-even point in dollars = Fixed expenses / Contribution margin ratio = 220500 / 50% 441000
Req 2 :
Increase in Contribution margin ( Increase in sales * Contribution margin ratio = 86000 * 50% ) 43000
(-) Increase in advertising budget 6700
Increase (decrease) in net operating income 36300
Increase in monthly net operating income 36300
Req 3 :
Revised selling price = Current selling price * ( 1 - % reduction ) = 30 * ( 1 - 10% ) 27
Current unit variable cost = Variable expenses / Units sold = 198000 / 13200 15
Revised fixed costs = Current fixed costs + Increase in advertising expense = 220500 + 33000 253500
Revised sales units = Current sales units * 2 = 13200 * 2 26400
Sales ( 26400 * 27 ) 712800
(-) Variable expenses ( 26400 * 15 ) 396000
Contribution margin 316800
(-) Fixed expenses 253500
Net operating income (loss) 63300
Req 4 :
Revised unit variable cost = Current unit variable cost + 0.50 = 15 + 0.50 15.50
Units sales to attain target profit = ( Target profit + Fixed expenses ) / ( Selling price - Unit variable cost ) = ( 4900 + 220500 ) / ( 30 - 15.50 ) 15545
Req 5A :
Current unit variable cost = Variable expenses / Units sold = 198000 / 13200 15
Revised unit variable cost = 15 - 3 12
Revised fixed expenses = 220500 + 50000 270500
Contribution margin per unit = Selling price - Unit variable cost = 30 - 12 18
CM ratio = Contribution margin per unit / Selling price = 18 / 30 60%
Break-even point in unit sales = Fixed costs / Contribution margin per unit = 270500 / 18 15028
Break even point in dollar sales = Fixed costs / CM ratio = 270500 / 60% 450833
Req 5B :
Not automated Automated
Total Per unit % Total Per unit %
Sales 618000 30 100% 618000 30 100%
Variable expenses 309000 15 50% 247200 12 40%
Contribution margin 309000 15 50% 370800 18 60%
Fixed expenses 220500 270500
Net operating income 88500 100300
Req 5C :
Answer : Yes
Explanation : If the company sells 20600 units, then the company will earn more net operating income if it automates its operations.